What mass of water will be produced from 2.70 moles of Ca(OH)2 reacting with HCl?
1
Expert's answer
2020-11-22T11:16:45-0500

The equation of Ca(OH)2 reacting with HCl is

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

Clearly,

1 mole of Ca(OH)2 reacting with HCl to produce 2 mole of H2O

2.70 moles of Ca(OH)2 reacting with HCl to produce (2×2.70) mole of H2O

= 5.4 moles of H2O

Molar mass of H2O = 18 g/mole

So,

mass of 5.4 moles of H2O = (18×5.4) g

= 97.2 g


Hence,

97.2 g of water will be produced from 2.70 moles of Ca(OH)2 reacting with HCl.